# Galaxy Zoo: Cosmic Dawn – morphological classifications for over 41 000 galaxies in the Euclid Deep Field North from the Hawaii Two-0 Cosmic Dawn survey

### v1

@@ 4.1.2 Aggregated classifications
In agreement with the H20 Cosmic Dawn team, the results of aggregated votes presented in this work, alongside the released classifications and subject images, are limited to the 45 742 subjects with either $z_{\rm phot}<2.5$ or for which the photometric redshift estimation failed, such that follow-up work on sources with higher estimated redshifts (3.4 per cent of the data set) can be carried out by the H20 Cosmic Dawn team themselves. For this redshift-limited sample, the cumulative number counts for each of the first responses – ‘Smooth’, ‘Features or Disk’, and ‘Star, Artifact, or Bad Zoom’ (referred to here as ‘Problem’) – are given as a function of photometric redshift in Figure 8. Limited to those with photometric redshifts as per the figure, the total number of subjects voted by more than 50 per cent of volunteers as being ‘Smooth’, ‘Features or Disk’, or ‘Problem’ subjects were 23 533, 2920, and 2346, respectively, while the remaining 7282 subjects’ votes were less certain. Including the subjects without photometric redshifts, these figures become 28 605 (63 per cent), 3201 (7 per cent), and 4123 (9 per cent), respectively, with 9297 (21 per cent) subjects’ votes less certain. As such, by separating out the ‘Problem’ subjects, this indicates the subject set contains at least 41 000 cutouts with central galaxies. For the sample of subjects with $z_{\rm phot}\gtrsim 1.0$, the majority of galaxies are seeing-limited so observationally biased towards being classified as ‘Smooth’. However, there are still extended objects in this sample, with 14 per cent of subjects classified as ‘Features or Disk’.
@@
Figure 10: Example GZCD subjects voted as being ‘Star, Artifact, or Bad Zoom’ subjects in response to the first question, having received the highest volunteer vote fractions for that response. Subjects are limited to those with at least ten volunteer votes. The hazy green and blue objects in the third, fourth and sixth images correspond to the ejected material surrounding the Cat’s Eye Nebula.
Some examples of classified subjects from the full data set are provided in Figures 9 and 10, which show subjects with the highest volunteer vote fractions for their response options. The first contains subjects with the highest vote fractions for ‘Features or Disk’, one of the responses to the first question presented to volunteers, with the majority of subjects with high vote fractions for this response containing clear spiral galaxies. Figure 10 on the other hand demonstrates an array of ‘Star, Artifact, or Bad Zoom’ subjects in response to the first question, including those featuring the most common image artifacts, which are then further classified into separate types of artifact in a follow-up question. From visual inspection of subjects voted as ‘Bad Image Zoom’, almost all were either zoomed-in portions of a nearby galaxy, or very zoomed out to the point of having no notable central galaxy. Regardless, these subjects did not undergo reprocessing as any galaxies in those cutouts were also already present as correct central galaxies in other cutouts. While over a thousand subjects were classified as stars in this project (see Section 5), only several tens of subjects were classified as artifacts. The vast majority of these were scattered light or saturation features, including the ejected material from the Cat’s Eye Nebula (the hazy blue-green objects in Figure 10), with no subjects receiving high vote fractions for the other responses, indicating the image reduction pipeline is robust against such artifacts. Nevertheless, subjects like these, especially the identified population of stars, can be used to improve the automated image reduction pipeline to reduce their presence in future data sets, such as acting as a training set for deep learning models.
More examples of classified subjects can be found in Appendix A for subjects classified as ‘Edge-On Disk’, ‘Major Disturbance’, ‘Clumps’, and ‘Ring’ (Figures 16 to 19, respectively). See Section 5 for the number counts of subjects for each question and response; more detailed investigations of the classified subjects, such as for identified ‘clumpy’ galaxies, are left for future work.
4.1.3 Tags, EROs and gravitational lenses
@@
In addition to aggregating volunteer votes for the decision tree of questions, another source of classifications comes from the tags that volunteers have assigned subjects on the Galaxy Zoo Talk boards - see Section 3.2. In that section it was highlighted that EROs should be tagged appropriately: across the objects below $z_{\rm phot}<2.5$, 2484 subjects were tagged as ‘#bright-red’ or ‘#extremely-red’; 789 tagged as ‘#dim-red’ or ‘#dark-red’; 572 tagged as ‘#bright-red-companion’; and 445 tagged as ‘#dim-red-companion’. However, these samples were found to be highly impure due to the subjective nature of defining colours, with many having a more orange appearance akin to many of the subjects in this data set. Hence, alternative algorithm-based methods may prove to be more effective in this regard, although the ‘#bright-red-companion’ tags provided a slightly purer sample: examples of which are shown in Figure 20 of Appendix A, which contain extremely red point-source objects away from the cutout centre. Elsewhere, a variety of tags were used to highlight objects volunteers found particularly beautiful or interesting, which were typically spiral, irregular or disturbed galaxies, as well as other rare objects - see Figure 21 in Appendix A for some examples.
@@
As discussed in Sections 3.1 and 3.5.2, we also aimed to search this data set for gravitational lenses, including through a dedicated Galaxy Zoo Mobile workflow set up to find lenses in subjects retired early by Zoobot. Classifications from the main Galaxy Zoo workflow were searched with a selection cut of at least 6 votes for ‘Lens or arc’, which from visual inspection led to a purer and equally complete sample compared to using cuts based on vote fractions. These 179 potential candidates, and all potential candidates from the Talk boards (467) and Galaxy Zoo Mobile (501), went through initial visual inspection vetting by the lead author to rule out clear false positives. Across the entire project, 122 unique lens candidates were found as a result, with some duplicates obtained by different means: 32 from the main Galaxy Zoo workflow; 56 tagged by volunteers in the Talk boards, plus another 5 not tagged as lenses, but as ‘Rings’; and 63 from Galaxy Zoo Mobile. These were then classified by a team of experts in the Euclid Consortium using both HSC and Euclid imaging, narrowing the final list of lenses to the following 57: 15 new grade A lenses, 9 grade B lenses, 27 grade C lenses, and another 6 lenses already known to Euclid (all grade A). The 57 lenses are shown in Table 1 and Figure 11, and are listed in the main catalogue of Euclid Collaboration: Walmsley et al. (2025b). Of these, 23 (2), 34 (11), and 23 (21) lenses were found (uniquely) through the main workflow, Talk tags, and Galaxy Zoo Mobile, respectively. 23 (40 per cent) were found by more than one method, with the vast majority of overlap coming from the main workflow and Talk tags (the latter finding all 23). Hence, most of the lenses were found through Talk tags, which proved to be more complete than the main workflow albeit with a higher rate of false positives, while the dedicated Galaxy Zoo Mobile workflow found by far the most candidates not seen by other methods. This highlights the importance of using multiple approaches to find rare objects where completeness is of great importance.
@@ 4.2 Zoobot classifications
Figure 14: Confusion matrices of subject classifications from Zoobot and volunteers for each question in the decision tree. Matrices are limited to subjects with at least 40 volunteer votes, and further limited to those with ‘high confidence’ classifications (i.e. those with volunteer vote fractions >0.8) and, for a given question, at least 5 volunteer votes to reduce the ‘ground truth’ uncertainty.
Figure 12 shows the mean absolute deviations between Zoobot’s predicted vote fractions and those of volunteers. Across the possible response options, 43 per cent, 68 per cent, and 86 per cent had mean deviations below 0.1, 0.2 and 0.3, respectively. These deviations are typically larger than the $\sim$10 per cent seen in previous iterations (Walmsley et al., 2022a; Euclid Collaboration: Walmsley et al., 2025a) due to smaller sources from deeper imaging and increased confusion noise from background galaxies. Some of the more accurately predicted responses are for the initial ‘smooth-or-featured’ question, the ‘how-rounded’ question for ‘Smooth’-voted galaxies, and the ‘merging’ question, which received the largest number of volunteer votes due to the structure of the decision tree (Figure 4) combined with the majority of subjects being classed as ‘Smooth’ in this subject set (see Figure 8). Also, despite the ‘problem’ and ‘artifact’ questions being new additions to this Galaxy Zoo iteration, most of their responses were still predicted within 10 per cent of the volunteer’s results. The largest deviations came from perhaps less clear-cut and more subjective response options, such as the shapes of edge-on bulges, the tightness of spiral arm winding, the ‘Can’t Tell’ option for spiral arm count and ‘Other’ for artifact type: volunteers’ responses for these can vary more widely while Zoobot may choose more specific responses based on hard-to-spot features.
To analyse these differences further, the mean absolute deviations are shown again in Figure 13, now binned as functions of redshift. As can be seen, there are large uncertainties in the binned deviations between Zoobot and volunteer vote fractions, however the deviations do increase or decrease with redshift for several question responses, such as the increase for all ‘disk-edge-on’ and ‘clumps’ responses and the decrease for all ‘problem’ responses. Interestingly, for the majority of responses with the highest deviations in Figure 12, there is very little dependence on redshift, suggesting these deviations are affected primarily by limited training of Zoobot on this survey data, rather than an unavoidable limitation of the model (or volunteer accuracy) when applied to higher-redshift subjects.
@@
Confusion matrices for each question are presented in Figure 14, comparing the numbers of subjects classified as each response by volunteers and by Zoobot. Corresponding classification metrics are given in Table 2, and as with the other comparisons in this section, these matrices and metrics are limited to subjects with at least 40 volunteer responses (i.e. ‘fully classified’ by volunteers). In accordance with the above, we only consider subjects that have received at least 5 total votes in response to a given question, and here they are further limited to those with ‘high confidence’ classifications (defined as those with volunteer vote fractions >0.8) as presented in Euclid Collaboration: Walmsley et al. (2025a): Figure 13 of that work highlights that Zoobot prediction errors tend to occur primarily when a subject’s volunteer vote fraction lies within the range 0.4-0.6, with Zoobot becoming increasingly accurate for subjects with a more confident volunteer vote fraction (closer to 0 or 1). Even so, for the GZCD data set, excluding subjects with ‘uncertain’ labels produced only a minor effect on Zoobot accuracy. It is clear from the results that Zoobot performed exceptionally across the majority of questions and responses, even predicting with 100 per cent accuracy how volunteers would vote for responses to the first question (‘smooth-or-featured’). It should be noted however that this is fairly expected behaviour for a well-trained model, as it is testing here on the same data set used to train it (albeit only one subset of the training data). The model primarily struggled on accurately predicting volunteer votes for ‘Boxy’ edge-on bulges, ‘Weak’ bars, ‘Medium’ spiral winding, ‘Large’ bulge sizes, and those undergoing ‘Major Disturbances’, however it is clear from Figure 14 that in this subject set the ground truth sets for these responses were very small.
[…]
4.3 Quantitative Tests of our Qualitative Classifications
@@ 5 Data Release
Alongside this paper, we release catalogues containing the aggregated classifications made by volunteers and by Zoobot for GZCD, along with the set of PNG/JPG cutout images used in the project and the associated metadata for each subject, all available from Zenodo (Pearson et al., 2025). We present an overview here, and additional documentation can be found alongside the data itself. All subjects (both cutouts and tabular data) are identified by their H20 IDs, and, in addition to the data mentioned in Section 2.4, the metadata includes the Galaxy Zoo Talk board tags assigned to each subject as well as API links to various search databases. As discussed in Section 4.1.2, the data set consists of 45 742 subjects with either photometric redshifts $z_{\rm phot}<2.5$ or for which the photometric redshift estimation failed. The cutouts are split into two groups: the 15 745 PNG subjects fully classified by volunteers prior to Zoobot active learning, and the 29 997 JPG subjects classified with the help of active learning. Hence, many of those in the latter set will have less than 40 votes recorded in the classification catalogues.
@@
Table 3: Example suggested selection cuts exploring the aggregated vote fractions $f$, and the number counts $N_{\rm sample}$ of subjects meeting those cuts for classifications made by volunteers and by Zoobot. Number counts for volunteer classifications are limited to subjects with at least 10 volunteer votes. See Table 2 of Euclid Collaboration: Walmsley et al. (2025a) for more suggested cuts, which broadly apply to this data set as well. The table here includes suggested cuts for questions and responses not included in such previous works, with $f_{\rm problem}$ representing vote fractions for the ‘Star, Artifact, or Bad Zoom’ response. For ‘Lens or Arc’, ‘Irregular’ and ‘Ring’, no number counts are given for Zoobot as it was not trained to predict responses to the ‘rare features’ question.
@@
Regardless of the catalogue, these vote fractions can be used to place selection cuts so as to interrogate the subject set. For example, one may use $f_{\rm smooth}>0.7$ (where $f_{\rm smooth}$ corresponds to the header ‘smooth-or-featured_smooth’) to select a fairly pure sample of galaxies classed as ‘Smooth’. Such selection cuts can also be combined in order to correctly interpret classifications made further along the branches of the decision tree: one might extract edge-on disks by specifying $f_{\rm featured}>0.3$ and $f_{\rm yes\ edge-on}>0.8$ (regarding headers ‘smooth-or-featured_featured-or-disk_fraction’ and ‘disk-edge-on_yes’, respectively). It should be noted that there is no perfect cut for all users, as the balance of purity and completeness will depend on your science goal: a strict cut produces a purer sample, say, for follow-up observations, while a more relaxed cut would produce a sample whose distribution can be weighted by the subjects’ vote fractions. To help obtain fairly pure samples, some examples of approximate selection cuts are provided in Table 3 based on visual inspection, and more can be found in Walmsley et al. (2022a). In particular, the table provides cuts for questions and responses not previously included in recent data releases, including ‘Artifact’, ‘Clumps’ and some ‘rare features’ responses, although for all cuts the reader should take care to examine the purity of their samples and adjust the cuts accordingly. For each of these selection cuts, the table also provides the corresponding number counts of subjects as classified by volunteers and by Zoobot.
@@
It is clear from the table that for the first question, many galaxies have not been confidently classified, that is, have not received high vote fractions for any of the three responses. This highlights that these votes are entirely based on visual morphology rather than true morphology, and hence subjects can be limited by image resolution and confusion noise. As mentioned in Sections 3.5.1 and 4.1.2, the vote fractions have not been debiased with regard to redshift and so the majority of subjects with $z_{\rm phot}\gtrsim 1.0$ are seeing-limited and hence likely classified as ‘Smooth’. With the GZCD data set spread across a fairly wide redshift range, it is also possible that many of the galaxies are indeed less clearly classifiable due to unusual features or in-progress morphological evolution. It should be noted that, given the PSF and lack of clear diffraction spikes for all but the brightest of stars in the deep HSC imaging, one area of difficulty in this Galaxy Zoo iteration came from faint stars being selected as subjects in the cutout-creation process, which made it challenging for volunteers (and hence a retrained Zoobot) to distinguish between stars and smooth spherical galaxies. As a result, the catalogue includes a large number of subjects classified as stars, which may well be PSF-limited stars or smooth galaxies matching their appearance: we caution the reader to take care when utilising the data set for such objects, and highlight that a stricter cut of $f_{\rm smooth}>0.8$ for example tends toward more elliptical-shaped galaxies.

### v2

@@ 4.1.2 Aggregated classifications
In agreement with the H20 Cosmic Dawn team, the results of aggregated votes presented in this work, alongside the released classifications and cutout images, are limited to the 45 742 subjects with either $z_{\rm phot}<2.5$ or for which the photometric redshift estimation failed, such that follow-up work on sources with higher estimated redshifts (3.4 per cent of the data set) can be carried out by the H20 Cosmic Dawn team themselves. For this redshift-limited sample, the cumulative number counts for each of the first responses – ‘Smooth’, ‘Features or Disk’, and ‘Star, Artifact, or Bad Zoom’ (referred to here as ‘Problem’) – are given as a function of photometric redshift in Figure 8. Limited to those with photometric redshifts as per the figure, the total number of subjects voted by more than 50 per cent of volunteers as being ‘Smooth’, ‘Features or Disk’, or ‘Problem’ subjects were 23 533, 2920, and 2346, respectively, while the remaining 7282 subjects’ votes were less certain. Including the objects without photometric redshifts, these figures become 28 605 (63 per cent), 3201 (7 per cent), and 4123 (9 per cent), respectively, with 9297 (21 per cent) subjects’ votes less certain. As such, by separating out the ‘Problem’ images, this indicates the data set contains at least 41 000 cutouts with central galaxies. For the sample of subjects with $z_{\rm phot}\gtrsim 1.0$, the majority of galaxies are seeing-limited so observationally biased towards being classified as ‘Smooth’. However, there are still extended objects in this sample, with 14 per cent of subjects classified as ‘Features or Disk’.
@@
Figure 10: Example GZCD subjects voted as being ‘Star, Artifact, or Bad Zoom’ in response to the first question, having received the highest volunteer vote fractions for that response. Subjects are limited to those with at least ten volunteer votes. The hazy green and blue objects in the third, fourth and sixth images correspond to the ejected material surrounding the Cat’s Eye Nebula.
Some examples of classified subjects from the full data set are provided in Figures 9 and 10, which show those with the highest volunteer vote fractions for their response options. The first contains subjects with the highest vote fractions for ‘Features or Disk’, one of the responses to the first question presented to volunteers, with the majority of those with high vote fractions for this response containing clear spiral galaxies. Figure 10 on the other hand demonstrates an array of ‘Star, Artifact, or Bad Zoom’ subjects in response to the first question, including those featuring the most common image artifacts, which are then further classified into separate types of artifact in a follow-up question. From visual inspection of those voted as ‘Bad Image Zoom’, almost all were either zoomed-in portions of a nearby galaxy, or very zoomed out to the point of having no notable central galaxy. Regardless, these cutouts did not undergo reprocessing as any galaxies in them were also already present as correct central galaxies in other cutouts. While over a thousand subjects were classified as stars in this project (see Section 5), only several tens were classified as artifacts. The vast majority of these were scattered light or saturation features, including the ejected material from the Cat’s Eye Nebula (the hazy blue-green objects in Figure 10), with none receiving high vote fractions for the other responses, indicating the image reduction pipeline is robust against such artifacts. Nevertheless, subjects like these, especially the identified population of stars, can be used to improve the automated image reduction pipeline to reduce their presence in future data sets, such as acting as a training set for deep learning models.
More examples of classified subjects can be found in Appendix A for those classified as ‘Edge-On Disk’, ‘Major Disturbance’, ‘Clumps’, and ‘Ring’ (Figures 16 to 19, respectively). See Section 5 for the number counts of subjects for each question and response; more detailed investigations of the classified objects, such as for identified ‘clumpy’ galaxies, are left for future work.
@@ 4.1.3 Tags, EROs and gravitational lenses
In addition to aggregating volunteer votes for the decision tree of questions, another source of classifications comes from the tags that volunteers have assigned subjects on the Galaxy Zoo Talk boards - see Section 3.2. In that section it was highlighted that EROs should be tagged appropriately: across the objects below $z_{\rm phot}<2.5$, 2484 were tagged as ‘#bright-red’ or ‘#extremely-red’; 789 tagged as ‘#dim-red’ or ‘#dark-red’; 572 tagged as ‘#bright-red-companion’; and 445 tagged as ‘#dim-red-companion’. However, these samples were found to be highly impure due to the subjective nature of defining colours, with many having a more orange appearance akin to many of the objects in this data set. Hence, alternative algorithm-based methods may prove to be more effective in this regard, although the ‘#bright-red-companion’ tags provided a slightly purer sample: examples of which are shown in Figure 20 of Appendix A, which contain extremely red point-source objects away from the cutout centre. Elsewhere, a variety of tags were used to highlight objects volunteers found particularly beautiful or interesting, which were typically spiral, irregular or disturbed galaxies, as well as other rare objects - see Figure 21 in Appendix A for some examples.
@@
As discussed in Sections 3.1 and 3.5.2, we also aimed to search this data set for gravitational lenses, including through a dedicated Galaxy Zoo Mobile workflow set up to find lenses in subjects retired early by Zoobot. Classifications from the main Galaxy Zoo workflow were searched with a selection cut of at least 6 votes for ‘Lens or arc’, which from visual inspection led to a purer and equally complete sample compared to using cuts based on vote fractions. These 179 potential candidates, and all potential candidates from the Talk boards (467) and Galaxy Zoo Mobile (501), went through initial visual inspection vetting by the lead author to rule out clear false positives. Across the entire project, 122 unique lens candidates were found as a result, with some duplicates obtained by different means: 32 from the main Galaxy Zoo workflow; 56 tagged by volunteers in the Talk boards, plus another 5 not tagged as lenses, but as ‘Rings’; and 63 from Galaxy Zoo Mobile. These were then classified by a team of experts in the Euclid Consortium’s Strong Lensing Science Working Group using both HSC and Euclid imaging (Euclid Collaboration, 2025b), narrowing the final list of lenses to the following 57: 15 new grade A lenses, 9 grade B lenses, 27 grade C lenses, and another 6 lenses already known to Euclid (all grade A). The 57 lenses are shown in Table 1 and Figure 11, and are listed in the main catalogue of the above publication. Of these, 23 (2), 34 (11), and 23 (21) lenses were found (uniquely) through the main workflow, Talk tags, and Galaxy Zoo Mobile, respectively. 23 (40 per cent) were found by more than one method, with the vast majority of overlap coming from the main workflow and Talk tags (the latter finding all 23). Hence, most of the lenses were found through Talk tags, which proved to be more complete than the main workflow albeit with a higher rate of false positives, while the dedicated Galaxy Zoo Mobile workflow found by far the most candidates not seen by other methods. This highlights the importance of using multiple approaches to find rare objects where completeness is of great importance.
@@ 4.2 Zoobot classifications
Figure 14: Confusion matrices of classifications from Zoobot and volunteers for each question in the decision tree. Matrices are limited to subjects with at least 40 volunteer votes, and further limited to those with ‘high confidence’ classifications (i.e. those with volunteer vote fractions >0.8) and, for a given question, at least 5 volunteer votes to reduce the ‘ground truth’ uncertainty.
Figure 12 shows the mean absolute deviations between Zoobot’s predicted vote fractions and those of volunteers. Across the possible response options, 43 per cent, 68 per cent, and 86 per cent had mean deviations below 0.1, 0.2 and 0.3, respectively. These deviations are typically larger than the $\sim$10 per cent seen in previous iterations (Walmsley et al., 2022a; Euclid Collaboration, 2025a) due to smaller sources from deeper imaging and increased confusion noise from background galaxies. Some of the more accurately predicted responses are for the initial ‘smooth-or-featured’ question, the ‘how-rounded’ question for ‘Smooth’-voted galaxies, and the ‘merging’ question, which received the largest number of volunteer votes due to the structure of the decision tree (Figure 4) combined with the majority of subjects being classed as ‘Smooth’ in this data set (see Figure 8). Also, despite the ‘problem’ and ‘artifact’ questions being new additions to this Galaxy Zoo iteration, most of their responses were still predicted within 10 per cent of the volunteer’s results. The largest deviations came from perhaps less clear-cut and more subjective response options, such as the shapes of edge-on bulges, the tightness of spiral arm winding, the ‘Can’t Tell’ option for spiral arm count and ‘Other’ for artifact type: volunteers’ responses for these can vary more widely while Zoobot may choose more specific responses based on hard-to-spot features.
To analyse these differences further, the mean absolute deviations are shown again in Figure 13, now binned as functions of redshift. As can be seen, there are large uncertainties in the binned deviations between Zoobot and volunteer vote fractions, however the deviations do increase or decrease with redshift for several question responses, such as the increase for all ‘disk-edge-on’ and ‘clumps’ responses and the decrease for all ‘problem’ responses. Interestingly, for the majority of responses with the highest deviations in Figure 12, there is very little dependence on redshift, suggesting these deviations are affected primarily by limited training of Zoobot on this survey data, rather than an unavoidable limitation of the model (or volunteer accuracy) when applied to higher-redshift objects.
@@
Confusion matrices for each question are presented in Figure 14, comparing the numbers of subjects classified as each response by volunteers and by Zoobot. Corresponding classification metrics are given in Table 2, and as with the other comparisons in this section, these matrices and metrics are limited to subjects with at least 40 volunteer responses (i.e. ‘fully classified’ by volunteers). In accordance with the above, we only consider subjects that have received at least 5 total votes in response to a given question, and here they are further limited to those with ‘high confidence’ classifications (defined as those with volunteer vote fractions >0.8) as presented in Euclid Collaboration (2025a): Figure 13 of that work highlights that Zoobot prediction errors tend to occur primarily when a subject’s volunteer vote fraction lies within the range 0.4-0.6, with Zoobot becoming increasingly accurate for those with a more confident volunteer vote fraction (closer to 0 or 1). Even so, for the GZCD data set, excluding subjects with ‘uncertain’ labels produced only a minor effect on Zoobot accuracy. It is clear from the results that Zoobot performed exceptionally across the majority of questions and responses, even predicting with 100 per cent accuracy how volunteers would vote for responses to the first question (‘smooth-or-featured’). It should be noted however that this is fairly expected behaviour for a well-trained model, as it is testing here on the same data set used to train it (albeit only one subset of the training data). The model primarily struggled on accurately predicting volunteer votes for ‘Boxy’ edge-on bulges, ‘Weak’ bars, ‘Medium’ spiral winding, ‘Large’ bulge sizes, and those undergoing ‘Major Disturbances’, however it is clear from Figure 14 that in this subject set the ground truth sets for these responses were very small.
[…]
4.3 Quantitative Tests of our Qualitative Classifications
@@ 5 Data Release
Alongside this paper, we release catalogues containing the aggregated classifications made by volunteers and by Zoobot for GZCD, along with the set of PNG/JPG cutout images used in the project and the associated metadata for each subject, all available from Zenodo (Pearson et al., 2025). We present an overview here, and additional documentation can be found alongside the data itself. All subjects (both cutouts and tabular data) are identified by their H20 IDs, and, in addition to the data mentioned in Section 2.4, the metadata includes the Galaxy Zoo Talk board tags assigned to each subject as well as API links to various search databases. As discussed in Section 4.1.2, the data set consists of 45 742 subjects with either photometric redshifts $z_{\rm phot}<2.5$ or for which the photometric redshift estimation failed. The cutouts are split into two groups: the 15 745 PNGs fully classified by volunteers prior to Zoobot active learning, and the 29 997 JPGs classified with the help of active learning. Hence, many of those in the latter set will have less than 40 votes recorded in the classification catalogues.
@@
Table 3: Example suggested selection cuts exploring the aggregated vote fractions $f$, and the number counts $N_{\rm sample}$ of subjects meeting those cuts for classifications made by volunteers and by Zoobot. Number counts for volunteer classifications are limited to subjects with at least 10 volunteer votes. See Table 2 of Euclid Collaboration (2025a) for more suggested cuts, which broadly apply to this data set as well. The table here includes suggested cuts for questions and responses not included in such previous works, with $f_{\rm problem}$ representing vote fractions for the ‘Star, Artifact, or Bad Zoom’ response. For ‘Lens or Arc’, ‘Irregular’ and ‘Ring’, no number counts are given for Zoobot as it was not trained to predict responses to the ‘rare features’ question.
@@
Regardless of the catalogue, these vote fractions can be used to place selection cuts so as to interrogate the data set. For example, one may use $f_{\rm smooth}>0.7$ (where $f_{\rm smooth}$ corresponds to the header ‘smooth-or-featured_smooth’) to select a fairly pure sample of galaxies classed as ‘Smooth’. Such selection cuts can also be combined in order to correctly interpret classifications made further along the branches of the decision tree: one might extract edge-on disks by specifying $f_{\rm featured}>0.3$ and $f_{\rm yes\ edge-on}>0.8$ (regarding headers ‘smooth-or-featured_featured-or-disk_fraction’ and ‘disk-edge-on_yes’, respectively). It should be noted that there is no perfect cut for all users, as the balance of purity and completeness will depend on your science goal: a strict cut produces a purer sample, say, for follow-up observations, while a more relaxed cut would produce a sample whose distribution can be weighted by the subjects’ vote fractions. To help obtain fairly pure samples, some examples of approximate selection cuts are provided in Table 3 based on visual inspection, and more can be found in Walmsley et al. (2022a). In particular, the table provides cuts for questions and responses not previously included in recent data releases, including ‘Artifact’, ‘Clumps’ and some ‘rare features’ responses, although for all cuts the reader should take care to examine the purity of their samples and adjust the cuts accordingly. For each of these selection cuts, the table also provides the corresponding number counts of subjects as classified by volunteers and by Zoobot.
@@
It is clear from the table that for the first question, many galaxies have not been confidently classified, that is, have not received high vote fractions for any of the three responses. This highlights that these votes are entirely based on visual morphology rather than true morphology, and hence limited by image resolution and confusion noise. As mentioned in Sections 3.5.1 and 4.1.2, the vote fractions have not been debiased with regard to redshift and so the majority of objects with $z_{\rm phot}\gtrsim 1.0$ are seeing-limited and hence likely classified as ‘Smooth’. With the GZCD data set spread across a fairly wide redshift range, it is also possible that many of the galaxies are indeed less clearly classifiable due to unusual features or in-progress morphological evolution. It should be noted that, given the PSF and lack of clear diffraction spikes for all but the brightest of stars in the deep HSC imaging, one area of difficulty in this Galaxy Zoo iteration came from faint stars being selected as subjects in the cutout-creation process, which made it challenging for volunteers (and hence a retrained Zoobot) to distinguish between stars and smooth spherical galaxies. As a result, the catalogue includes a large number of objects classified as stars, which may well be PSF-limited stars or smooth galaxies matching their appearance: we caution the reader to take care when utilising the data set for such objects, and highlight that a stricter cut of $f_{\rm smooth}>0.8$ for example tends toward more elliptical-shaped galaxies.